Using the method of undetermined coefficients, find the general solution of the differential equation y^(iv) -2y^''' +y^''=3 e^-x+2 e^x x + e^-x sin (x ) 


1
Expert's answer
2021-04-15T06:57:13-0400

"y^{(4)}-2y^{'''}+y''=3e^{-x}+2xe^x+e^{-x}sinx"


"m^4-2m^3+m^2=0"

"m^2(m^2-2m+1)=0"

"m^2(m-1)^2=0"

"m_1=m_2=0,\ m_3=m_4=1"

The complementary solution:

"y_c=c_1+c_2x+c_3e^x+c_4xe^x"


Solve for particular solutions, using the method of undetermined coefficients:

"y_{p1}=Ae^{-x}"

"Ae^{-x}+2Ae^{-x}+Ae^{-x}=3e^{-x}"

"4A=3\implies A=3/4"

"y_{p1}=\frac{3e^{-x}}{4}"


"y_{p2}=(Ax^3+Bx^2)e^x"

"y'_{p2}=(3Ax^2+2Bx)e^x+(Ax^3+Bx^2)e^x=(Ax^3+(3A+B)x^2+2Bx)e^x"


"y''_{p2}=(3Ax^2+2(3A+B)x+2B)e^x+(Ax^3+(3A+B)x^2+2Bx)e^x="

"=(Ax^3+(6A+B)x^2+2(3A+2B)x+2B)e^x"


"y'''_{p2}=(3Ax^2+2(6A+B)x+2(3A+2B))e^x+(Ax^3+(6A+B)x^2+"

"+2(3A+2B)x+2B)e^x=(Ax^3+(9A+B)x^2+(18A+6B)x+6A+6B)e^x"

"y^{(4)}_{p2}=(3Ax^2+2(9A+B)x+18A+6B)e^x+(Ax^3+(9A+B)x^2+"

"+(18A+6B)x+6A+6B)e^x=(Ax^3+(12A+B)x^2+(36A+8B)x+"

"+24A+12B)e^x"

"36A+8B-2(18A+6B)+6A+4B=2"

"6A=2\implies A=1/3"

"24A+12B-2(6A+6B)+2B=0"

"12A+2B=0\implies B=-2"

"y_{p2}=(\frac{x^3}{3}-2x^2)e^x"


"y_{p3}=e^{-x}(Acosx+Bsinx)"

"y'_{p3}=(Bcosx-Asinx)e^{-x}-(Acosx+Bsinx)e^{-x}=((B-A)cosx-"

"-(A+B)sinx)e^{-x}"

"y''_{p3}=((A-B)sinx-(A+B)cosx)e^{-x}-((B-A)cosx-(A+B)sinx)e^{-x}="

"=(2Asinx-2Bcosx)e^{-x}"

"y'''_{p3}=(2Acosx+2Bsinx)e^{-x}-(2Asinx-2Bcosx)e^{-x}="

"=((2A+2B)cosx+(2B-2A)sinx)e^{-x}"

"y^{(4)}_{p3}=((2B-2A)cosx-(2A+2B)sinx)e^{-x}-((2A+2B)cosx+"

"+(2B-2A)sinx)e^{-x}=(-4Acosx-4Bsinx)e^{-x}"

"-4A-2(2A+2B)-2B=0"

"-8A-6B=0"

"-4B-2(2B-2A)+2A=1"

"-8B+6A=1"

"A=-3B\/4"

"-8B-18B/4=1\implies -50B=4\implies B=-2/25"

"A=3\/50"

"y^{(4)}_{p3}=e^{-x}(\frac{3cosx}{50}-\frac{2sinx}{25})"


The general solution:

"y=y_c+y_p"


"y(x)=c_1+c_2x+c_3e^x+c_4xe^x+\frac{3e^{-x}}{4}+(\frac{x^3}{3}-2x^2)e^x+e^{-x}(\frac{3cosx}{50}-\frac{2sinx}{25})"